OpenStudy (anonymous):

Factor trinomial below. Please show your work and check your answer. x^2 – 8x + 15

OpenStudy (lgbasallote):

so what's your answer?

OpenStudy (anonymous):

x=-b+root(b^2-4ac)/2*a x=8+root(64-60)/2 x=8+2/2 x=10/2 x=5 x=8-2/2 x=6/2 x=3

OpenStudy (goformit100):